# Language Model Council: Democratically Benchmarking Foundation Models on Highly Subjective Tasks

### v2

@@ Abstract
As Large Language Models (LLMs) have become more advanced, the search for efficient and accurate evaluation methods is ongoing. Many recent evaluations use other LLMs as judges to score outputs from other LLMs, often relying on a single large model like GPT-4o. However, using a single LLM judge is prone to intra-model bias, and many tasks – such as those related to emotional intelligence, creative writing, and persuasiveness – may be too subjective for a single model to judge fairly. We introduce the Language Model Council (LMC), where a group of LLMs collaborate to create tests, respond to them, and evaluate each other’s responses in a democratic fashion. Unlike previous approaches that focus on reducing cost or bias by using a panel of smaller models, our work examines the benefits and nuances of a fully inclusive LLM evaluation system. In a detailed case study on emotional intelligence, we deploy a council of 20 recent LLMs to rank each other on open-ended responses to interpersonal conflicts. Our results show that the LMC produces rankings that are more separable and more robust, and through a user study, we show that they are more consistent with human evaluations than any individual LLM judge. Using all LLMs for judging can be costly, however, so we use Monte Carlo simulations and hand-curated sub-councils to study hypothetical council compositions and discuss the value of the incremental LLM judge.
@@
Justin Zhao1, Flor Miriam Plaza-del-Arco2, Benjie Genchel1, Amanda Cercas Curry3 1Independent, 2Bocconi University, 3CENTAI justinxzhao@gmail.com, flor.plaza@unibocconi.it benjiegenchel@gmail.com, amanda.cercas@centai.eu https://llm-council.com
@@ 1 Introduction
Figure 1: Overview of the Language Model Council (LMC) evaluation framework. Using the same LLMs for test set formulation, task completion, and judging offers an equitable way to achieve a ranking by consensus on highly subjective tasks.
@@
Despite recent advances in Large Language Models (LLMs), evaluating their outputs remains an open challenge. Human evaluations are challenging, time-consuming and expensive, motivating the search for automatic metrics Novikova et al. (2017); Lowe et al. (2017) and evaluation setups (e.g. Zheng et al., 2024; Li et al., 2024b; Kocmi and Federmann, 2023; Shen et al., 2023; Kim et al., 2024). Conventional model evaluations use closed-ended questions that can be checked automatically such as MMLU Hendrycks et al. (2020). These static benchmarks are at risk of contamination Ravaut et al. (2024), prompting leaderboards to keep tests private or update questions periodically (e.g. Scale AI, 2024; White et al., 2024). However, performance on these questions is often misaligned with human preferences in real-world, open-ended contexts Chiang et al. (2024). For such tasks, automatic metrics like BLEU Papineni et al. (2002), ROUGE Lin (2004), or BLEURT Sellam et al. (2020) are common. These require reference responses and fail to reflect human preferences beyond a quality threshold Freitag et al. (2020). Arena-based methods enable reference-free evaluation by judging LLMs in head-to-head comparisons. The outcome of a battle between the responses of two LLMs can be evaluated based on objective outcomes Bianchi et al. (2024), strong model judges like GPT-4 Dubois et al. (2024b); Li et al. (2024a), or human judges Li et al. (2024a), with each outcome contributing to metrics like win rates or ELO scores Chiang et al. (2024).
@@
Strong LLM judges like GPT-4 can match both controlled and crowdsourced human preferences well, achieving over 80% agreement, the same level of agreement between humans, in a general-domain, arena-based setting Zheng et al. (2024). Unfortunately, it has also been observed that evaluator models tend to have their own biases, for example recognizing and preferring their own outputs over those of other models Dubois et al. (2024a); Panickssery et al. (2024), leading to the development of committees of LLM judges (such as Verga et al., 2024; Zhao et al., 2024).
@@
In a landscape of diverse LLM judges, how should we aggregate each model’s vote and determine a definitive ranking among them? This paper aims to contribute a unique perspective in the pursuit of a fully decentralized approach to evaluating LLMs. When ranking LLMs on competencies that may be too subjective for a single model to judge fairly, we explore the benefits of a fully inclusive evaluation network where each model contributes equally.
@@
Inspired by democratic organizations in human society, we introduce the Language Model Council (LMC), a framework for collective evaluation. The LMC operates in three stages: (1) collaborative test set creation by council members, (2) response generation by each LLM, and (3) evaluation of responses by the council as a collective jury.
@@
    We define and analyze key measures of LLM judging dynamics including: separability, pairwise positional consistency, agreement, and affinity, for the largest ensemble of LLM judges to date (to our knowledge).
@@
    Using all LLMs for judging can be costly, so we use Monte Carlo simulations and hand-crafted sub-councils to discuss the value of larger evaluation ensembles and the value of the incremental judge.
@@ 2 Related Work
Recently, subjective tasks and label variability in humans has received increased attention in the NLP community Alm (2011); Basile (2022); Plank (2022). LLMs also exhibit variability in their outputs as they inherit inconsistencies and biases from human data Hosking et al. (2023); Song et al. (2024); Bai et al. (2024a); Plaza-del-Arco et al. (2024); Koo et al. (2023). While LLM judges might not replicate human disagreement patterns exactly Lee et al. (2023); Dong et al. (2024), their differences in judgment are increasingly viewed as reflecting valid dissent. [^1] Notably, when humans disagreed with GPT-4, they considered its judgments reasonable in 75% of cases and sometimes revised their own answers Zheng et al. (2024).
LLM evaluation ensembles. Language-Model-as-an-Examiner Bai et al. (2024b) uses LLMs to interact with candidates through follow-up fact-oriented queries in the knowledge domain. Auto Arena Zhao et al. (2024) proposes an LLM committee to judge competitive multi-turn interactions between LLMs. PRD Li et al. (2023) allows LLMs to discuss evaluations and assigns higher voting weights based on ability. PRE Chu et al. (2024) selects a small group of reviewers to produce individual evaluations, then aggregates these evaluations through a chair model. DRPE Wu et al. (2023) uses multi-roleplayer prompting to simulate different roles with one base model, integrating them as votes for the final results. PoLL Verga et al. (2024) enhances cost-effectiveness by replacing one large judge with multiple smaller judges.
We build upon existing research by 1) focusing on a highly subjective case study on emotional intelligence where human agreement is inherently low, 2) emphasizing full inclusivity, where each LLM plays an equal role in determining the final rankings, and 3) engaging a large ensemble of diverse LLMs to study judging dynamics in greater depth. To our knowledge, this is the largest ensemble of LLM judges studied to date.
@@ 3 Case Study: Using the LMC to Rank LLMs on Emotional Intelligence
The LMC framework consists of three stages: 1) test set formulation, 2) response gathering, and 3) collective judging (Figure 1). These steps are run sequentially and are automated with LLM-based agents. [^2] We use “agents” to refer to LLMs acting simultaneously on sequential and systematic evaluation procedures. See Appendix I for all prompts.
@@ 3.1 Council member selection
Our selection of LLM council members was guided by several key considerations, including their widespread adoption within the AI community, availability of technical reports, well-supported API access, and performance on benchmarks like MMLU Hendrycks et al. (2020) and Chatbot Arena Chiang et al. (2024). We ensure a broad variety of LLMs by including models from eight different organizations and four countries, with a mix of open-source and closed-source models, small and large (Table 3 in Appendix A).
@@ 3.2 Test set formulation
To create a compelling open-ended test set for Emotional Application (EA), we build upon the EmoBench dataset, a publicly available, hand-crafted, theory-based English dataset designed for EA assessment Sabour et al. (2024). EmoBench consists of 200 emotionally balanced, handcrafted scenarios, e.g., “Sarah found out that her younger brother is being bullied at school, but he begged her not to tell their parents.” We use the LMC to transform the concise, closed-ended scenarios in EmoBench into richly described, open-ended dilemmas in the first person (Figure 27). Each of the 20 council members expands five scenarios, resulting in a dataset of 100 dilemmas, similar in scale to MT-Bench (80 questions). We manually review all expansions for EA suitability. [^3] An alternative procedure is to have all council members suggest expansions for each scenario and select the best through voting. This is feasible, but given the generally high quality of expansions, the added value is limited compared to simply adding new scenarios.
Relying on a single LLM to generate the entire test set, even a top performer like GPT-4o may introduce bias and limit perspectives. In a survey of 10 human respondents, 51% of preferred expansions were not written by GPT-4o. [^4] Human respondents were asked to choose, in a series of pairwise comparisons, which expanded dilemma would be better for an emotional intelligence test. Each comparison was between a response from GPT-4 and one from a randomly chosen council member. See Appendix E for more details. Inclusively constructing test examples also mitigates the risk of any single LLM’s generative idiosyncrasies AI4Science and Quantum (2023) from dominating the test pool.
@@ 3.3 Response gathering
After generating 100 dilemmas, each council member responds to every dilemma, yielding 2,000 total responses. To standardize response lengths across council members and minimize length bias in evaluation, the prompt suggests a 250-word limit (Figure 28). Responses exceeding the limit are truncated at the nearest sentence within the limit. [^5] Sentence splits are based on standard English end punctuation (. ! ?), as the experiment is conducted in English. Despite the suggested word limit, some council members consistently generated shorter responses (Figure 1). These responses are left unchanged.
3.4 LLM-as-a-Judge evaluation settings
Arena-style pairwise comparisons with a single reference model. We adopt the pairwise comparison setup of Chatbot Arena where responses are compared in a pairwise fashion in a series of battles (Figure 30). LLM rankings are based on expected win rates derived from an ELO scoring system Bai et al. (2022); Boubdir et al. (2023). We also use Bradley-Terry (BT) coefficients (Bradley and Terry, 1952) for better statistical estimation. We use a single reference model, Qwen-1.5-32B, for all pairwise battles. For details on how the reference model was chosen, refer to Appendix D. Following Li et al. (2024a); Chiang et al. (2024), we derive confidence intervals with 100 rounds of bootstrapping.
@@ 3.4 LLM-as-a-Judge evaluation settings
We query all LLMs with a temperature of 0 and with granular comparison options without ties (A>>B, A>B, B>A, B>>A). The reasoning behind these choices is detailed in Appendix C. We use Chain-of-Thought (CoT) prompting Wei et al. (2022) to generate discussion before giving judgments.
@@ 3.5 Characterizing LLM judges
We leverage the many-to-many interactions between LLMs to define key judging qualities of LLMs in an ensemble setting.
Separability measures how confidently models can be distinguished in the final rankings. We adopt this metric from Li et al. (2024a), which calculates separability as the percentage of model pairs with non-overlapping confidence intervals. Higher separability means better differentiation between models. The success of score separation depends on three factors: judge discrimination, test discrimination, and participant abilities. They all contribute, and any one of them could reduce the separation to 0. In the LMC framework, potentially all three factors are governed by stochastic LLMs that are non-deterministic even with temperature=0 Chann (2023). This makes it challenging to pinpoint which component would provide the most improvement if separability is low. Our analysis of separability focuses on LLM judging, as the test set, participants, and their responses remain constant across different LMC configurations.
Pairwise Positional Consistency (PPC) measures how often a judge gives consistent results when the order of the two model responses in a pairwise comparison is swapped (Appendix C.2). For instance, if the judge ranks A>B in one comparison and B>A when the positions are swapped (a battle couplet), the judge is considered "consistent" because the preference remained the same independent of position. Position bias measures how much the LLM judge favors a specific position (either the first or second) and we define it to be $1-ppc$. On the 4-point preference scale, a battle couplet is still considered consistent as long as the relative ranking remains consistent overall — fine-grained differences such as (A>>B, B>A) or (B>>A, A>B) are tolerated. Table 10 lists all possible couplets and their consistency mappings. Conviction is the percentage of strong votes (A>>B or B>>A).
@@ 3.6 Human study
To validate the LMC’s evaluations, we conduct a human study mirroring that of the LMC: human raters are asked to choose the best of two responses to a given dilemma. Note that the goal is to replicate human preferences in terms of general ranking, not to model the distribution of preferences.
Due to budget constraints, we select nine LLM council members from our pool of 20 to be rated for this study (listed in Figure 3) with variety in terms of model size, open- versus closed-source, and company origin in mind. We recruit participants via crowdsourcing on Prolific. [^9] https://www.prolific.com/ We recruited a total of 102 participants. Each response were rated by 11 participants on average, with a total of 1343 ratings. Detailed information on our recruitment process, quality control, and participant demographics is provided in Appendix E.

### v4

@@ Abstract
As Large Language Models (LLMs) continue to evolve, evaluating them remains a persistent challenge. Many recent evaluations use LLMs as judges to score outputs from other LLMs, often relying on a single large model like GPT-4o. However, using a single LLM judge is prone to intra-model bias, and many tasks – such as those related to emotional intelligence, creative writing, and persuasiveness – may be too subjective for a single model to judge fairly. We introduce the Language Model Council (LMC), where a group of LLMs collaborate to create tests, respond to them, and evaluate each other’s responses to produce a ranking in a democratic fashion. Unlike previous approaches that focus on reducing cost or bias by using a panel of smaller models, our work examines the benefits and nuances of a fully inclusive LLM evaluation system. In a detailed case study on emotional intelligence, we deploy a council of 20 recent LLMs to rank each other on open-ended responses to interpersonal conflicts. Our results show that the LMC produces rankings that are more separable and more robust, and through a user study, we show that they are more consistent with human evaluations than any individual LLM judge. Using all LLMs for judging can be costly, however, so we use Monte Carlo simulations and hand-curated sub-councils to study hypothetical council compositions and discuss the value of the incremental LLM judge.
@@
Justin Zhao1, Flor Miriam Plaza-del-Arco2, Benjamin Genchel1, Amanda Cercas Curry3 1Independent, 2Bocconi University, 3CENTAI Institute https://llm-council.com
@@ 1 Introduction
Figure 1: Overview of the Language Model Council (LMC) evaluation framework. By using the same LLMs for test set formulation, task completion, and judging, the framework offers an equitable way to achieve an inclusive, consensus-based ranking.
@@
As Large Language Models (LLMs) continue to advance, evaluating their outputs remains a significant challenge. Manual human evaluations are time-consuming and expensive, motivating the need for automatic metrics Novikova et al. (2017); Lowe et al. (2017) and evaluation methods (e.g. Zheng et al., 2023; Li et al., 2024c; Kocmi and Federmann, 2023; Shen et al., 2023). Conventional model evaluations rely on closed-ended questions that can be checked automatically such as MMLU Li et al. (2024a). However, these static benchmarks are vulnerable to data contamination Ravaut et al. (2024) and are often misaligned with human preferences in real-world, open-ended contexts Chiang et al. (2024). For evaluating open-ended responses, automatic metrics like BLEU Papineni et al. (2002), ROUGE Lin (2004), and BLEURT Sellam et al. (2020) are common. These require reference responses that may be expensive to collect and yet may still fail to reflect human preferences beyond a quality threshold Freitag et al. (2020).
Arena-based methods enable reference-free evaluation by comparing LLMs in head-to-head matchups. The outcome of a battle between the responses of two LLMs can be evaluated based on objective measures Bianchi et al. (2024), ratings from strong model judges like GPT-4 Dubois et al. (2023); Li et al. (2024b), or human judges Chiang et al. (2024), with each outcome contributing to metrics like win rates or ELO scores Li et al. (2024b).
@@
Strong LLM judges like GPT-4 can match both controlled and crowdsourced human preferences well, achieving over 80% agreement, the same level of agreement between humans, in a general-domain, arena-based setting Zheng et al. (2023). Unfortunately, it has also been observed that evaluator models tend to have their own biases, for example recognizing and preferring their own outputs over those of other models Dubois et al. (2024); Panickssery et al. (2024), which can be mitigated by using a committee of LLM judges (e.g. Verga et al., 2024; Zhao et al., 2024).
@@
In a landscape of diverse LLM judges, can we account for each model’s opinions and still establish a definitive ranking amongst them? This paper aims to contribute a unique perspective in the pursuit of a fully decentralized approach to evaluating LLMs. When ranking LLMs on competencies that may be too subjective for a single model to judge fairly, we investigate the dynamics and benefits of an inclusive evaluation network where all models contribute equally to the ranking process.
@@
Inspired by democratic organizations in human society, we introduce the Language Model Council (LMC), a framework for collective evaluation for a group of LLMs. The LMC operates in three stages: (1) a test set is formulated with equal participation from all council members, (2) the test set is administered to all council members to complete, and (3) the responses are evaluated by the council as a collective jury.
@@
    We define and analyze key measures of LLM judging dynamics including: separability, pairwise positional consistency, agreement, and affinity, for the largest ensemble of LLM judges to date.
@@
    We use Monte Carlo simulations and hand-crafted sub-councils to discuss the value of larger evaluation networks and the value of the incremental judge.
We release all data, code, leaderboard, and web demo at https://llm-council.com.
@@ 2 Related Work
Subjective tasks and label variability in humans have received increased attention in the NLP community Alm (2011); Basile (2022); Plank (2022). LLMs also exhibit variability in their outputs as they inherit inconsistencies and biases from human data Hosking et al. (2024); Song et al. (2024); Plaza-del-Arco et al. (2024); Bai et al. (2024); Plaza-del-Arco et al. (2024); Koo et al. (2024). While LLM judges might not replicate human disagreement patterns exactly Lee et al. (2023); Dong et al. (2024), their differences in judgment are increasingly viewed as reflecting valid dissent. Notably, when humans disagreed with GPT-4, they considered its judgments reasonable in 75% of cases and sometimes revised their own answers Zheng et al. (2023).
LLM evaluation ensembles. Recent work explores using LLMs as evaluators through structured interactions, committees, weighted voting, and role-playing techniques. Language-Model-as-an-Examiner Bai et al. (2023) uses LLMs to interact with candidates through follow-up fact-oriented queries in the knowledge domain. Auto Arena Zhao et al. (2024) proposes an LLM committee to judge competitive multi-turn interactions between LLMs. PRD Li et al. (2023) allows LLMs to discuss evaluations and assigns higher voting weights based on ability. PRE Chu et al. (2024) selects a small group of reviewers to produce individual evaluations, then aggregates these evaluations through a chair model. DRPE Wu et al. (2023) uses multi-roleplayer prompting to simulate different roles with one base model, integrating them as votes for the final results. PoLL Verga et al. (2024) enhances cost-effectiveness by replacing one large judge with multiple smaller judges.
We build upon existing research by (1) focusing on a highly subjective case study on emotional intelligence where human agreement is inherently low, (2) emphasizing full inclusivity, where each LLM plays an equal role in determining the final rankings, and (3) engaging a large ensemble of diverse LLMs to study judging dynamics in greater depth. To our knowledge, this is the largest ensemble of LLM judges studied to date.
@@ 3 Case Study: Using the LMC to Rank LLMs on Emotional Intelligence
The LMC framework consists of three stages: (1) test set formulation, (2) response gathering, and (3) collective judging (Figure 1). While the LMC framework is broadly applicable to a wide range of open-ended tasks, this paper presents a focused study on a subjective task of applying emotional intelligence (EI) in interpersonal conflict resolution.
See Appendix H for all prompts used in this case study.
3.1 Why the EI domain?
Unlike objective benchmarks like in coding and math, emotional intelligence (EI) benchmarks are often designed with subjectivity in mind. For example, they often incorporate ratings from a survey of humans for a single ground truth answer Wang et al. (2023); Sabour et al. (2024) and, in the case of multiple-choice questions, enabling multiple correct answers by measuring cosine similarity against a weighted distribution of choices Wang et al. (2023); Paech (2023). This thematic emphasis on multiple valid viewpoints in current EI benchmarks resonates with the LMC framework, which is itself designed to incorporate multiple LLM perspectives throughout the evaluation development process.
@@ 3.2 Council member selection
Our selection of LLM council members was guided by several key considerations, including their widespread adoption within the AI community, availability of technical reports, well-supported API access, and performance on benchmarks like MMLU Li et al. (2024a) and Chatbot Arena Chiang et al. (2024). We ensure a broad variety of LLMs by including models from eight different organizations and four countries, with a mix of open-source and closed-source models, small and large (Table 8).
@@ 3.3 Test set formulation
To create a compelling, open-ended test set for EI, we build upon the EmoBench dataset, a publicly available, hand-crafted, theory-based English dataset designed for EI assessment Sabour et al. (2024). EmoBench consists of 200 emotionally balanced, handcrafted scenarios, e.g., “Sarah found out that her younger brother is being bullied at school, but he begged her not to tell their parents.” We solicit the council to expand EmoBench’s concise scenarios into richly described dilemmas in the first person (see Figure 28 for an example). Each of the 20 council members expands five scenarios, resulting in a test set of 100 dilemmas, similar in scale to MT-Bench (80 questions). We manually review all expansions for EI suitability. [^1] Our manual review resulted in no omissions, though some submitted expansions required minor edits to remove preambles like ’Here is the expanded dilemma…’.
Relying on a single LLM to generate the entire test set – even a top performer like GPT-4o – may introduce bias and limit perspectives. In a survey of 10 human respondents evaluating potential expansions for the test set, 51% of the preferred expansions were those not authored by GPT-4o. [^2] Human respondents were asked to choose, in a series of pairwise comparisons, which expanded dilemma would be better for an emotional intelligence test. Each comparison was between a response from GPT-4 and one from a randomly chosen council member (Figure 21). Inclusively constructing test examples also mitigates the risk of any single LLM’s generative idiosyncrasies AI4Science and Quantum (2023) from dominating the test pool.
An alternative approach would be to have all council members propose expansions for each scenario and select the best through voting. While this may be just as rigorous from a democratic perspective, given the generally high quality of expansions, we opted to use a balanced set of submitted expansions directly.
@@ 3.4 Response gathering
After expanding 100 dilemmas, each council member responds to every dilemma, yielding 2,000 total responses. To standardize response lengths across council members and preemptively minimize length bias in evaluation, the prompt suggests a 250-word limit (Figure 29). Responses exceeding the limit are truncated at the nearest sentence within the limit. [^3] Sentence splits are based on standard English end punctuation (. ! ?), as the experiment is conducted in English. Despite the suggested word limit, some council members consistently generated shorter responses (Figure 1). These responses are left unchanged.
3.5 Collective judging
Arena-style pairwise comparisons with a single reference model. We adopt the pairwise comparison setup of Chatbot Arena where responses are compared in head-to-head matchups (see prompt in Figure 31). LLM rankings are determined by expected win rates using an ELO scoring system Bai et al. (2022); Boubdir et al. (2023), with Bradley-Terry (BT) coefficients (Bradley and Terry, 1952) applied for improved statistical estimation. Following Li et al. (2024b), confidence intervals are derived through 100 rounds of bootstrapping. Like Dubois et al. (2024); Li et al. (2024b), we use a single reference model for all pairwise battles. However, instead of GPT-4, we use Qwen-1.5-32B. For details on how the reference model was chosen, refer to Appendix C.
@@ 3.5 Collective judging
We query all LLMs with a temperature of 0 and with granular comparison options without ties (A>>B, A>B, B>A, B>>A). We use Chain-of-Thought (CoT) prompting Wei et al. (2022) to generate discussion before giving judgments. The reasoning behind all of these choices is detailed in Appendix B.
@@ 3.6 Characterizing LLM judges
We leverage the many-to-many interactions between LLMs to define key judging qualities of LLMs in an ensemble setting such as the LMC.
Separability measures how confidently models can be distinguished in the final rankings. We adopt this metric from Li et al. (2024b), which defines separability as the percentage of model pairs with non-overlapping confidence intervals, where higher separability indicates better differentiation. The success of score separation depends on three factors: judge discrimination, test discrimination, and participant abilities. They all contribute, and any one of them could reduce the separation to zero. In the LMC framework, all three are influenced by non-deterministic LLMs, which makes it challenging to isolate which factor contributes most to low separability, for example. Our analysis of separability focuses on LLMs as judges, as the test set, participants, and responses remain constant across different LMC configurations.
Pairwise Positional Consistency (PPC) measures how often a judge gives consistent results when the order of the two responses in a pairwise comparison is swapped. For instance, if the judge ranks A>B in one comparison and B>A when the positions are swapped (a rating couplet), the judge is considered "consistent" because the preference remained the same independent of position. Position bias measures how much the LLM judge favors a specific position (either the first or second) and we define it to be $1-ppc$. On the 4-point preference scale, a rating couplet is still considered consistent as long as the relative ranking remains consistent overall — fine-grained differences such as (A>>B, B>A) or (B>>A, A>B) are tolerated. Table 10 lists all possible couplets and their consistency mappings. Conviction is the raw percentage of strong votes (A>>B or B>>A).
@@ 3.7 Human study
To validate the LMC’s evaluations, we conduct a human study mirroring that of the LMC’s EI test. Human raters are asked to select the better response from a pair presented for each dilemma. The goal is to assess alignment with human preferences in overall ranking, rather than to model the exact distribution of preferences.
We select nine LLM council members from our pool of 20 to be rated for this study (Figure 3). [^7] The decision to evaluate only nine models was driven solely by budget constraints. Given the cost of human evaluations (€9/hour), the number of models studied had to be limited. However, these nine models were carefully chosen to ensure diversity in model size, openness, and company origin. Participants were recruited via crowdsourcing on Prolific. [^8] https://www.prolific.com/ A total of 102 participants took part in the study, with each response evaluated by an average of 11 raters, resulting in 1,343 total ratings. Further details on recruitment, quality control, and participant demographics are in Appendix D.